What can be the sale of a strangle?

Strangles: Long a CE and a PE on the same K and Maturity
Bottom Vertical Combination: Purchasing Strangle
K1P = $50 K2C = $60
PP= $5 CP= $3
Stock Price Long a K2Call Long a K1Put Profits
$20
$25
$30
$35
$40
$45
$50
$55
$60
$65
$70
$75
$80
$85
$90
$95
Answer #1

Long call profit = Max(St - X, 0) - Call price

Long put profit = Max(X - St, 0) - Put price

Profits = Long call profit + Long put profit

Can you please upvote? Thank You :-)

B KIP Long Strangle $50 K2C $5 CP $60 $3 PP $10 $5 Stock Price Long call profit Long put profit $20 ($3) $25 $25 ($3) $20 $30

Screenshot with formulas

B 50 Long Strangle K2C CP KIP PP 60 Stock Price Profits 20 11 12 13 14 Long call profit =MAX(B7-$E$3,0) - $E$4 =MAX(B8-$E$3,0